OpenStudy (horsegirl27):

Pleeeeeeeeeeeeaaaaaassssseeeeeeeeeeeee help me!!!!!!!!! medal and fan!!!!!!!!!!!!!!!!!!!!!

OpenStudy (horsegirl27):

http://prntscr.com/6v8t5i

OpenStudy (horsegirl27):

Part A: The tables does represent a relation that is a function because none of the x values repeat. Part B: f(h)=15h+5 f(8)=15(8)+5 f(8)=120+5 f(8)=125 here's my work, I just need part C... can someone help?

OpenStudy (anonymous):

hi

OpenStudy (horsegirl27):

hello, can you help me?

OpenStudy (anonymous):

haha sure

OpenStudy (horsegirl27):

Ok! Please explain and help me out, I feel like once you explain, I'll know, but right now I'm lost! XD

OpenStudy (anonymous):

k wat u need help with :)

OpenStudy (horsegirl27):

Part C of the question I posted at the top :) I posted my work for parts A and B too, so if there's any mistakes correct me, and then I believe u need it to answer the question.

OpenStudy (anonymous):

o hehe im silly

OpenStudy (horsegirl27):

haha, it's fine :P

OpenStudy (anonymous):

give me a sec

OpenStudy (horsegirl27):

Ok, take your time :)

OpenStudy (anonymous):

lets seee

OpenStudy (anonymous):

do i do part b or just c

OpenStudy (horsegirl27):

Just C, under the question, I posted my work for A and B so u could see :)

OpenStudy (anonymous):

o im so weird i was doin b sowwy let me do c haha

OpenStudy (horsegirl27):

That's ok haha

OpenStudy (anonymous):

yay now me lost ugh wait wat grade u in

OpenStudy (horsegirl27):

8th

OpenStudy (anonymous):

i gtg fan me and ill help later k bye buddy hehe

OpenStudy (horsegirl27):

ok lol

OpenStudy (horsegirl27):

@rational @rvc @nincompoop @tkhunny please help me :)

OpenStudy (anonymous):

sowwy i gtg go i really wanted to help bbye

OpenStudy (horsegirl27):

Ok, I'll be sure to tag you next time I need help though!

OpenStudy (rational):

part a and part b look good

OpenStudy (horsegirl27):

Yay! Can you help with C?

OpenStudy (rational):

part c : f(h) = 15h + 5 f(r) = 15r + 5 so the ordered pair would be (r, 15r+5)

OpenStudy (horsegirl27):

Oh, ok, thank you!!

OpenStudy (rational):

yw

OpenStudy (horsegirl27):

Can you help with another?

OpenStudy (rational):

wil try, post..

OpenStudy (horsegirl27):

Ok. http://prntscr.com/6v9497

OpenStudy (horsegirl27):

I don't have any work for this one

OpenStudy (rational):

for part a, you need to figure out the slope of each function

OpenStudy (horsegirl27):

Ok, how do I do that when it's in a table like this?

OpenStudy (rational):

look at the table of f(x) pick any two points and see if you can find the slope

OpenStudy (horsegirl27):

And how do i do it when written out?

OpenStudy (horsegirl27):

as an equation

OpenStudy (rational):

may be pick the points from last two rows : (0, -1) (1, 9) slope= ?

OpenStudy (horsegirl27):

idk... :/

OpenStudy (rational):

use slope formula

OpenStudy (horsegirl27):

oh, ok one sec

OpenStudy (horsegirl27):

10/1??

OpenStudy (rational):

Yes! so the slope of f(x) is 10

OpenStudy (horsegirl27):

oh, ok!

OpenStudy (rational):

next find the slope of other function

OpenStudy (horsegirl27):

How do I do that?

OpenStudy (rational):

\[y=5x+1\] the given equation is in slope intercept form, so slope = \(5\)

OpenStudy (horsegirl27):

ooooooooohhhh

OpenStudy (horsegirl27):

So the first has a greater slope..

OpenStudy (rational):

Yep!

OpenStudy (horsegirl27):

ok, now for B...

OpenStudy (horsegirl27):

Now how do I find the y-intercept?

OpenStudy (rational):

y intercept is the value of function when x = 0 look at the table, whats the value of y when x = 0 ?

OpenStudy (horsegirl27):

-1?

OpenStudy (rational):

Correct!

OpenStudy (horsegirl27):

Yay, and the second has a y-intercept of 1, right?

OpenStudy (rational):

Yes! good job

OpenStudy (horsegirl27):

Thanks sooo much!! One more? It's the last one I need help with :)

OpenStudy (rational):

okay

OpenStudy (horsegirl27):

http://prntscr.com/6v9da9

OpenStudy (rational):

notice that the linear equation represents a line and the two points are given by (2, 225) (5, 480)

OpenStudy (horsegirl27):

ok

OpenStudy (rational):

start by finding the slope

OpenStudy (horsegirl27):

I got m=255/3

OpenStudy (rational):

simplify

OpenStudy (horsegirl27):

85?

OpenStudy (rational):

Yes slope, m = 85 pick any point and write the equation in point slope form

OpenStudy (horsegirl27):

is y-y1 y2-y1? same for x-x1?

OpenStudy (rational):

you have two points, pick one point first

OpenStudy (horsegirl27):

ok

OpenStudy (rational):

lets pick (2, 225) x1 = 2 y1 = 225 so the equation is y-225 = 85(x-2)

OpenStudy (rational):

rearragnge and put it in standard form

OpenStudy (horsegirl27):

oh, ok

OpenStudy (rational):

y-225 = 85(x-2) y-225 = 85x-170 85x-y = -55

OpenStudy (rational):

thats the required standard form

OpenStudy (horsegirl27):

oh yeah Ax-By=C, is how i learned it

OpenStudy (rational):

Yep

OpenStudy (horsegirl27):

So what do I do in part B?

OpenStudy (rational):

put it in slope intercept form

OpenStudy (horsegirl27):

oh, ok

OpenStudy (rational):

85x-y = -55 y = 85x + 55

OpenStudy (rational):

in function notation you get f(x) = 85x + 55

OpenStudy (horsegirl27):

ok

OpenStudy (horsegirl27):

Now, idk what to say for part C

OpenStudy (rational):

for part c : the function can be graphed by plotting the given two points: (2, 225), (5, 480) and drawing a line through them. label x axis with number of days rented and y axis with cost of rent

OpenStudy (horsegirl27):

THANK YOU I WANT TO GIVE YOU 100 MEDALS, THANKS SO MUCH!!!!!!!!!!!

OpenStudy (rational):

yw:)

OpenStudy (horsegirl27):

:D